OpenStudy (anonymous):

Which of the following is a solution of x2 + 6x = -22?

OpenStudy (shadeyghost):

do you mean 2x?

OpenStudy (anonymous):

no, it is x2, as in 1x^2

OpenStudy (shadeyghost):

OH ok

OpenStudy (anonymous):

the answer isn't pretty, so you'll have to use the quadratic formula, do you know it?

OpenStudy (anonymous):

-b =/- sqrt\[-b \pm \sqrt{-b^2-4ac }\]

OpenStudy (anonymous):

divided by 2a

OpenStudy (anonymous):

@pking8786

OpenStudy (anonymous):

@sidsiddhartha @hackworthaustin123

OpenStudy (anonymous):

That's the one, and the coefficients are ax^2 + bx+c so \[-6\pm \sqrt{36-(4*1*22)}/2a \]

OpenStudy (anonymous):

I've just worked it out, and are you sure you've not got a sign wrong/ because the one you gave doesn't have a solution. had it been +22 (thus making x2+6x-22=0) then you would have some answers

OpenStudy (anonymous):

-3 ± i -3 ± i -6 ± i -6 ± i

OpenStudy (anonymous):

@micahm @midhun.madhu1987

OpenStudy (micahm):

Ok so we have to find 2 numbers that multiply to 5 and add to -6. Since 5 is prime, only factors are 1 and 5. Now we set the parentheses up: ( )( ) = 0 Since we know that its x^2, simply just put an x in front of both parentheses's. (x )(x ) = 0 Ok and since the two factors have to add to -6, only 2 negatives can do that and we know that the last 2 numbers are -5 and -1 now since it has to add to -6 and multiply to 5. (Negative + negative = negative still so that works. Negative * Negative = positive. That works.) So now you get: (x-5)(x-1) = 0 Solve for x now: x = 5 or 1

OpenStudy (anonymous):

@micahm its AskingAlex, just my old account, lost my password for the other one

OpenStudy (midhun.madhu1987):

 Substitute and find the value of x.